The angular velocity of flywheel decreases uniformly from 12000rev/min to 3000 rev/min in 8 sec. Find the angular acceleration and the number of revolutions made by the wheel in the 8 sec interval. How many more seconds are required for the wheel to come to rest?
Answer #1

Initial rotational speed of the flywheel = N1 = 12000 rev/min

Rotational speed of the flywheel after 8 sec = N2 = 3000 rev/min

Initial angular speed of the flywheel = \omega1

\omega1 = 2\piN1/60

\omega1 = 2\pi(12000)/60

\omega1 = 1256.637 rad/s

Angular speed of the flywheel after 8 sec= \omega2

\omega2 = 2\piN2/60

\omega2 = 2\pi(3000)/60

\omega2 = 314.16 rad/s

Angular acceleration of the flywheel = \alpha

Time period = t1 = 8 sec

\omega2 = \omega1 + \alphat1

314.16 = 1256.637 + \alpha(8)

\alpha = -117.81 rad/s2

Negative sign as it is deceleration.

Angle turned by the flywheel in the 8 sec = \theta

\omega22 = \omega12 + 2\alpha\theta

(314.16)2 = (1256.637)2 + 2(-117.81)\theta

\theta = 6283.168 rad

Number of revolutions made by the flywheel in the 8 sec = n

n = \theta/(2\pi)

n = 6283.168/(2\pi)

n = 1000 rev

Additional time required for the wheel to come to rest = t2

Final angular speed of the flywheel = \omega3 = 0 rad/s (comes to rest)

\omega3 = \omega2 + \alphat2

0 = 314.16 + (-117.81)t2

t2 = 2.67 sec

Angular acceleration of the flywheel = -117.81 rad/s2

Number of revolutions made by the wheel in 8 sec = 1000 rev

Additional time required for the wheel to come to rest = 2.67 sec